# GCA-SUN: A Gated Context-Aware Swin-UNet for Exemplar-Free Counting

### v1

@@ III-A Experimental Settings
The Swin-T blocks are pre-trained on ImageNet-22k [20], and other modules are randomly initialized. AdamW optimizer [21] is employed for training, with an initial learning rate of 0.003, a decay rate of 0.95 and a batch size of 16. The model is trained with a warm-up period of 50 epochs. The input image size is $384\times 384$. Data augmentation [8] is employed to facilitate efficient training. Experiments are conducted using two NVIDIA RTX A5000 GPUs.
III-B Comparison with State-of-the-Art Methods
@@
Comparison experiments are conducted on the FSC-147 dataset. The results are summarized in Table I. Following the practice in [12, 8], we report the errors on the test and validation sets. We have the following observations. 1) The proposed GCA-SUN outperforms all the compared methods regarding test errors, while performing slightly poorer than DAVE [17] regarding validation errors. The superior results demonstrate its effectiveness for EFC tasks. It outperforms not only the dedicated methods for solving EFC tasks such as RepRPN-C [11] and RCC [12], but also state-of-the-art models for CAC tasks. Compared to the second-best method, CounTR [8], the performance gain is 0.71 for of MAE and 14.68 for RMSE. 2) Although the GCA-SUN performs slightly poorer than DAVE [17] in terms of validation errors, it significantly outperforms DAVE in terms of test errors, i.e., a performance gain of 1.14 for MAE and 11.30 for RMSE. DAVE tends to overfit to the validation set, while generalizing poorly on the test set. In contrast, the GCA-SUN generalizes well on the novel test set with minimal errors.
Method Test Set Val Set
MAE RMSE MAE RMSE
FamNet [1] CVPR’21 32.27 131.46 32.15 98.75
LOCA [16] ICCV’23 16.22 103.96 17.43 54.96
DAVE [17] CVPR’24 15.14 103.49 15.54 52.67
CounTR [8] BMVC’22 14.71 106.87 18.07 71.84
RepRPN-C [11] ACCV’22 26.66 129.11 29.24 98.11
RCC [12] CVPR’23 17.12 104.53 17.49 58.81
Proposed GCA-SUN 14.00 92.19 16.06 53.04
We visually compare the density maps on the FSC-147 dataset. As shown in Fig. 2, our method can capture fine-grained details of objects. CounTR sometimes generates density maps that do not accurately distinguish between individual objects in the map, e.g., in the fourth column, CounTR can’t identify the far-way small fruit, while our method can.
@@
Figure 2: Visual comparisons to CounTR [8] on the FSC-147 dataset.
III-C Cross-Domain Evaluation on CARPK Dataset
@@
Following [22], we conduct a cross-domain evaluation, training the model on FSC-147 [1] and directly evaluating on CARPK [19], with results summarized in Table II. The results for all compared methods are reproduced from [12, 8, 16] under the same settings. Our model has shown superior cross-domain performance compared with other methods, achieving a performance gain of 0.56 on MAE and 0.61 on RMSE compared to the previous best-performing method CounTR. Compared to the earlier EFC model [12], the gains are even more significant, highlighting GCA-SUN’s superior generalization over all compared methods.
TABLE II: Comparison with other methods on the CARPK dataset [19].
Methods MAE RMSE
LOCA [16] ICCV’23 16.84 19.72
@@
RCC [12] CVPR’23 21.38 26.61
Proposed GCA-SUN 10.96 13.95
III-D Visualization of GCAM
We visualize the effects of the proposed GCAM in Fig. 3. Sub-figure (b) and (c) are obtained by projecting the density maps into two two-dimensional spaces, followed by kernel density estimation [23] to calculate the density distribution. The resulting images are then normalized and visualized to illustrate the model’s focus areas. Clearly after applying GCAM, a general decrease of density values in the background areas (sky) can be observed, with the foreground objects (birds) becoming more prominent. This indicates that the GCAM module effectively enhances the representation of foreground tokens while suppressing irrelevant ones.
Figure 3: Visualization of the effects of GCAM.
III-E Ablation Study
We conduct a set of comprehensive ablation studies on the three major modules of proposed method on the FSC-147 dataset [1]. The results are summarized in Table III. The GCAM module alone significantly decreases the MAE on the test set by 1.77 and on the validation set by 2.33, highlighting its capability to enhance feature selectivity crucial for complex scenes. Similarly, utilizing GEFS alone or GAFU alone also greatly reduces the errors in both test set and validation set, demonstrating the importance of the gate mechanism in highlighting the relevant features while suppressing irrelevant ones. The full integration of all three components produces the most substantial enhancement, reducing MAE by 2.83 on the test set and by 3.34 on the validation set. This underscores the effectiveness of their synergistic interaction and affirms the component’s design.
@@ III-E Ablation Study
16.83 88.41 19.40 71.58
15.06 91.68 17.07 53.22
15.78 88.21 18.26 57.30
15.82 84.64 17.92 59.61
14.47 89.16 16.40 54.18
14.86 86.85 17.72 57.27
14.66 94.47 17.59 57.85
14.00 92.19 16.06 53.04
IV Conclusion
The proposed GCA-SUN effectively tackles the problems of exemplar-free counting by using a Swin-UNet architecture to directly map the input image to the density map of countable objects. The proposed GCAM exploits the attention information among the tokens of repetitive objects through the self-similarity matrix, and suppresses the features of irrelevant objects through a gate mechanism. The gate mechanism is also incorporated into the GEFS module and the GAFU module, which highlight the features most relevant to countable objects while suppressing irrelevant ones. Our experiments on the FSC-147 and CARPK datasets demonstrate that GCA-SUN outperforms state-of-the-art methods, achieving superior performance in both intra-domain and cross-domain scenarios.

### v2

@@ IV-A Experimental Settings
The Swin-T blocks are pre-trained on ImageNet-22k [29], and other modules are randomly initialized. AdamW optimizer is employed, with an initial learning rate of 0.003, a decay rate of 0.95 and a batch size of 16. The model is trained with a warm-up period of 50 epochs and 300 epochs in total. The input image size is $384\times 384$. Data augmentation [9] is employed to facilitate efficient training. All experiments are conducted using two NVIDIA RTX A5000 GPUs.
IV-B Comparison Results on FSC-147 Dataset
@@
Comparison experiments are conducted on the FSC-147 dataset. The results are summarized in Table I. Following the practice in [13, 9], we report the errors on the test and validation sets. We have the following observations. 1) The proposed GCA-SUNet outperforms all the compared methods regarding test errors, while performing slightly poorer than DAVE [18] regarding validation errors. The superior results demonstrate its effectiveness for EFC tasks. It outperforms not only the dedicated methods for solving EFC tasks such as RepRPN-C [12] and RCC [13], but also state-of-the-art models for CAC tasks. Compared to the second-best method, CounTR [9], the performance gain is 0.63 for of MAE and 16.47 for RMSE. 2) Although the GCA-SUNet performs slightly poorer than DAVE [18] in terms of validation errors, it significantly outperforms DAVE in terms of test errors, i.e., a performance gain of 1.06 for MAE and 18.72 for RMSE. DAVE performs well on the validation set and generalizes poorly on the test set. In contrast, the GCA-SUNet generalizes well on the novel test set with minimal errors.
@@
Figure 2: Visual comparisons of the generated density maps with CounTR [9] and RCC [13] on the FSC-147 dataset.
We visually compare the density maps with CounTR [9] and RCC [13] on the FSC-147 dataset. As shown in Fig. 2, our method can capture fine-grained details of objects. In contrast, CounTR sometimes generates density maps that do not accurately distinguish between individual objects in the map, e.g., in the fourth column, CounTR can’t identify the far-way small fruit, while our method can.
IV-C Cross-Domain Evaluation
@@
Following [30], we conduct a cross-domain evaluation, i.e., training the model on the FSC-147 dataset [7] while directly evaluating on the CARPK dataset [19]. The results are summarized in Table II. The results for all the compared methods are obtained from [13, 9, 17] under the same settings. As shown in Table II, our model shows superior cross-domain performance compared with other methods, achieving a performance gain of 0.56 on MAE and 0.61 on RMSE compared to the previous best-performing method CounTR. Compared to the earlier EFC model [13], the gains are even more significant, highlighting GCA-SUNet’s superior generalization over the compared methods.
TABLE II: Training on the FSC-147 dataset [7] while testing on the CARPK dataset [19].
Category Method Test Set
MAE RMSE
CAC LOCA [17] ICCV’23 16.84 19.72
@@
EFC RCC [13] CVPR’23 21.38 26.61
Proposed GCA-SUNet 10.96 13.95
IV-D Ablation Study of Major Modules
We ablate the three major modules of proposed method on the FSC-147 dataset [7]. Swin-UNet [24] serves as the baseline model. The results are summarized in Table III. The GCAM module alone significantly reduces the MAE by 2.32 on the test set and by 1.72 on the validation set compared with the baseline, underscoring its capability to enhance feature selectivity that is essential for complex scenes. Similarly, utilizing GEFS or GAFU individually also greatly reduces errors on both test and validation sets, showcasing the proposed gate mechanism’s ability to highlight relevant features while suppressing irrelevant ones. The combination of these modules also achieves significant performance gains. Specifically, the overall framework with all three components achieves the most substantial improvement compared with the baseline, reducing the MAE by 4.78 on the test set and by 5.83 on the validation set. The results show the effectiveness of the synergistic interaction among different modules and emphasize the individual contribution of each major component.
@@ IV-D Ablation Study of Major Modules
18.86 89.00 22.85 85.16
16.54 85.20 21.13 81.30
16.63 80.33 20.13 62.98
16.87 84.36 21.60 79.37
15.27 84.81 20.08 81.58
15.07 78.79 18.58 59.17
15.21 85.71 17.97 55.67
14.08 84.77 17.02 55.37
IV-E Analysis of GCAM
Figure 3: Feature maps before/after the $\rm 1^{st}$ GCAM in the Swin-T enconder.
As shown in Table III, GCAM contributes most significantly to the overall performance gain. We further visualize its effects as the density maps in Fig. 3. Following the practice in [13], sub-figure (b) and (c) are obtained by projecting the density maps into two-dimensional spaces. Clearly by applying GCAM, the noisy estimation in the background areas (sky) is significantly reduced, with the foreground objects (birds) becoming more prominent. The results show that the proposed GCAM effectively enhances the representation of foreground tokens while suppressing irrelevant ones in the background.
Figure 4: Evaluation of GCAM in various object density scenarios.
We further compare the results with/without the GCAM across scenarios of different object densities in Fig. 4. For different object densities, the GCAM could consistently reduce the MAE, signifying its importance. On the other hand, Fig. 4 reveals a clear trend of increasing error as the number of objects grows. Indeed, as the density of objects increases, they become smaller and counting tasks become more challenging. Future work will focus on optimizing the proposed approach to improve its performance in such challenging scenes.
IV-F Model Complexity Analysis
Tab. IV compares the inference time, model size, FLOPs and training epochs with the previous best method, CounTR [9]. The proposed model requires inference time comparable to CounTR. It significantly outperforms CounTR in terms of accuracy, at a higher but reasonable computational cost.
TABLE IV: Comparison of inference time, model size and FLOPs, and training epochs.
Method Inf. time Parameters FLOPs Training
(s) (M) (G) (epochs)
CounTR [9] 0.08 100 84 1000
Proposed GCA-SUNet 0.11 192 173 300
V Conclusion
The proposed GCA-SUNet effectively tackles the problems of exemplar-free counting by using a Swin-UNet architecture to directly map the input image to the density map of countable objects. The proposed GCAM exploits the attention information among the tokens of repetitive objects through the self-similarity matrix, and suppresses the features of irrelevant objects through a gate mechanism. The gate mechanism is also incorporated into the GEFS module and the GAFU module, which highlight the features most relevant to countable objects while suppressing irrelevant ones. Our experiments on the FSC-147 and CARPK datasets demonstrate that GCA-SUNet outperforms state-of-the-art methods, achieving superior performance in both intra-domain and cross-domain scenarios.